If a train is moving toward you and at a speed of 50 m/sec and emits a whistle sound of pitch 1500 Hz, what will be the frequency with which you hear the pitch?
1
Expert's answer
2011-06-07T05:53:23-0400
Frequency f0, then an observer stationary relative to the medium detects waves with a frequency f given by
f = v/(v + vs) f0
where vs is positive if the source is moving away from the observer, and negative if the source is moving towards the observer.
In this case v is the speed of sound ( approximately 343 m/s)
Vs=-50 m/sec
So f=(v/(v+vs )) f0

f = (343/(343-50))*1500 ≈ 1755.97 Hz

Answer : f ≈ 1755.97 Hz